OpenStudy (anonymous):

x+5 ----- =15 (please solve and show work/explain 2

TheSmartOne (thesmartone):

Multiply both sides by 2 and finally subtract 5.

TheSmartOne (thesmartone):

So \[2(\frac{ x+5 }{ 2 })=15 \times 2\]

TheSmartOne (thesmartone):

Then \[x+5=30\]

TheSmartOne (thesmartone):

x=30-5 x=?